# Boosting Text-to-Chart Retrieval with Semantic Insights

### v2

@@ 6.5. Exp-4: The Generalizability of Synthesized Semantic Insights
Overall Results. As shown in Table  6, the introduction of semantic insights has brought significant performance improvements to different models. For the CLIP-DPR, in the precise query scenario, R@10 and MRR@10 increased by 4.62 and 4.18 percentage points, respectively; in the fuzzy query scenario, the improvement was more significant, with R@10 increasing by 8.39 percentage points. The UniVL-DR achieved greater performance gains after the introduction of semantic insights, especially in the fuzzy query scenario, with R@10 increasing by 10.69 percentage points and MRR@10 increasing by 8.82 percentage points. This result fully proves that the semantic insights we designed have good versatility and can effectively enhance the performance of retrieval models with different architectures, especially showing stronger robustness when processing fuzzy queries. This further verifies the universal value of multi-dimensional semantic insights for chart retrieval tasks.
Finding 5: We found that fine-tuning other models using semantic insights can also significantly improve their retrieval capabilities, proving the generalizability of the three semantic insights.
Table 6. The Impact of Training with Semantic Insights.
@@
R@10 MRR@10 R@10 MRR@10
CLIP-DPR (Liu et al., 2022) 20.00 8.90 10.69 6.60
w/ three semantic insights 24.62 13.08 19.08 10.97
$\uparrow$4.62 $\uparrow$4.18 $\uparrow$8.39 $\uparrow$4.37
UniVL-DR (Liu et al., 2022) 17.44 6.59 9.92 6.26
w/ three semantic insights 26.15 13.25 20.61 15.08
$\uparrow$8.71 $\uparrow$6.66 $\uparrow$10.69 $\uparrow$8.82
Table 7. Ablation Study on Center Crop. W/ direct resize means we directly resize the image to the required size instead of cropping the chart to the target size.
Models R@5 R@10 MRR@5 MRR@10 NDCG@10
Urban1k
Long-CLIP-B 94.1 97.4 85.3 85.8 88.6
w/ direct resize 93.2 96.1 84.7 85.1 87.8
$\downarrow$0.9 $\downarrow$1.3 $\downarrow$0.6 $\downarrow$0.7 $\downarrow$0.8
Long-CLIP-L 96.7 98.2 90.5 90.7 92.6
w/ direct resize 95.1 97.5 88.5 88.9 90.9
$\downarrow$1.6 $\downarrow$0.7 $\downarrow$2.0 $\downarrow$1.8 $\downarrow$1.5
Chart-to-Text
Long-CLIP-B 87.29 90.38 79.46 79.87 82.44
w/ direct resize 98.28 98.85 94.10 94.18 95.35
$\uparrow$10.99 $\uparrow$8.47 $\uparrow$14.64 $\uparrow$14.31 $\uparrow$12.91
Long-CLIP-L 88.51 91.03 80.87 81.20 83.61
w/ direct resize 99.35 99.78 97.00 97.06 97.75
$\uparrow$10.84 $\uparrow$8.75 $\uparrow$16.13 $\uparrow$15.86 $\uparrow$14.14
6.6. Exp-5: The Impact of Center Crop Strategy
Experiment Settings. We mentioned in Section 6.1 that for the text-to-chart retrieval task, we replaced the traditional center crop preprocessing method in the CLIP-based models with a direct resize strategy. To verify the necessity of this modification, we selected Long-CLIP-B and Long-CLIP-L as the base models and conducted comparative tests on the urban1k and Chart-to-Text datasets, respectively. Urban1k (Zhang et al., 2025) is a retrieval dataset containing natural images of urban scenes, which contains a variety of urban buildings, streets, and landscapes, representing a typical natural image retrieval task.
@@ 6.6. Exp-5: The Impact of Center Crop Strategy
Figure 6. Case study on Top-1 Text-to-Chart Retrieval Results of the ChartFinder and CLIP.
Overall Results. The results in Table  7 show that the preprocessing strategies have very different effects on different retrieval tasks. On the Urban1k dataset, direct resize caused a slight decrease in model performance, with Long-CLIP-B and Long-CLIP-L decreasing by 0.9 and 1.5 percentage points on average, respectively. However, on the Chart-to-Text dataset, the same strategy brought significant improvements, with the two models gaining approximately 12.3 and 13.1 percentage points in performance, respectively. This shows that in chart retrieval tasks, retaining the complete structure of the chart (including edge axis labels, legends, etc.) is more important than maintaining precise proportions, proving that our adjustments to chart-specific preprocessing strategies are necessary and effective.
Finding 6: In natural images, a lot of important semantic information is often in the center of the image, so the center crop processing allows the model to pay more attention to the query target. However, in the chart, important information such as the title, x-axis, and y-axis names are all around the chart, which makes it easy for the center crop to cut out important information.
6.7. Exp-6: Case Study
@@
7. Related Work
Text-to-Chart Retrieval. Previous work on text-to-chart chart retrieval has primarily focused on two approaches: visualization recommendation systems (Vartak et al., 2017; Ye et al., 2024; Qin et al., 2018b, a) and natural language interfaces (Shen et al., 2022; Tang et al., 2022; Luo et al., 2018b, 2022b, 2021). In the context of natural language interfaces, systems like Olio (Setlur et al., 2023) and BOLT (Srinivasan and Setlur, 2023) allow users to express their analytical intents or design exploration using natural language for chart retrieval. On the other hand, visualization recommendation systems such as SlopeSeeker (Bendeck et al., 2024) and Snowy (Srinivasan and Setlur, 2021) focus on guiding users in exploring and discovering interesting patterns and charts within the data. Although much progress has been made in these two approaches, they still struggle to understand the user’s query intent. We believe embedding the query and charts into the same latent space will shine a light on Text-to-Chart Retrieval. With the development of MLLM (Li et al., 2022, 2023; Yuan et al., 2021; Xiao et al., 2024b) models, we can see text-to-chart as a cross-modal retrieval task, which has been driven by large-scale image-text paired datasets and pre-trained models like CLIP (Radford et al., 2021). Several variants have enhanced CLIP’s performance, such as EVA-CLIP (Sun et al., 2023), which improves retrieval by using EVA (Fang et al., 2024) as the visual encoder, and Jina-CLIP-v2 (Xiao et al., 2024a; Koukounas et al., 2024) which enhances the text encoder (Sturua et al., 2024) for better retrieval. Other approaches, like UniVL-DR (Liu et al., 2022), MagicLens (Zhang et al., 2024), VISTA (Zhou et al., 2024a), and MARVEL (Zhou et al., 2024b), fine-tune CLIP on large, diverse datasets to improve model’s retrieval abilities. However, MLLM-based methods have limited research focused on chart retrieval. This paper addresses this gap by injecting semantic insights into ChartFinder based on contrastive learning.
Chart Captions with Semantic Insights. Studies show that rich captions can significantly enhance CLIP-based models’ performance (Li et al., 2024). For instance, VeCLIP (Lai et al., 2025) boosts performance by injecting extra visual information into captions, while DreamLIP (Zheng et al., 2025) achieves the same effect as CLIP 300M data using just one-tenth of the long caption data. Another approach to enrich semantic information is to use multiple captions for the same image. Llip (Lavoie et al., 2024) suggests that each image should have captions from various angles. VisText (Tang et al., 2023) introduces a dataset for statistical charts with two levels of semantic information per chart. We extend this idea by proposing three levels of semantic insights to reflect the semantic depth of charts. Moreover, while VisText’s manually annotated data is limited, we generate semantic insights using Llama 3.1 (Dubey et al., 2024) 70B based on metadata, ensuring both quality and quantity.
8. Conclusion
In this paper, we introduced ChartFinder, a novel text-to-chart retrieval model that combines contrastive learning with the CLIP architecture to align both visual and semantic features of charts. By leveraging synthesized semantic insights of charts, our model effectively retrieves relevant charts based on a user’s textual query. We also presented CRBench, the first benchmark for text-to-chart retrieval in real-world BI scenarios. Experiments show that ChartFinder outperforms existing methods, achieving significant improvements in both precise (up to 66.9% NDCG@10) and fuzzy query tasks on CRBench.

### v4

@@ A.3 Exp-4: The Generalizability of Synthesized Semantic Insights
Overall Results. As shown in Table 6, the introduction of semantic insights has brought significant performance improvements to different models. For the CLIP-DPR, in the precise query scenario, R@10 and MRR@10 increased by 4.62 and 4.18 percentage points, respectively; in the fuzzy query scenario, the improvement was more significant, with R@10 increasing by 8.39 percentage points. The UniVL-DR achieved greater performance gains after the introduction of semantic insights, especially in the fuzzy query scenario, with R@10 increasing by 10.69 percentage points and MRR@10 increasing by 8.82 percentage points. This result fully proves that the semantic insights we designed have good versatility and can effectively enhance the performance of retrieval models with different architectures, especially showing stronger robustness when processing fuzzy queries. This further verifies the universal value of multi-dimensional semantic insights for chart retrieval tasks.
Finding 4: We found that fine-tuning other models using semantic insights can also significantly improve their retrieval capabilities, proving the generalizability of the three semantic insights.
A.4 Exp-5: The Impact of Center Crop Strategy
Experiment Settings. We mentioned in Section 5 Training Details that for the text-to-chart retrieval task, we replaced the traditional center crop preprocessing method in the CLIP-based models with a direct resize strategy. To verify the necessity of this modification, we selected Long-CLIP-B and Long-CLIP-L as the base models and conducted comparative tests on the urban1k and Chart-to-Text datasets, respectively. Urban1k Zhang et al. (2025) is a retrieval dataset containing natural images of urban scenes, which contains a variety of urban buildings, streets, and landscapes, representing a typical natural image retrieval task.
Overall Results. The results in Table 7 show that the preprocessing strategies have very different effects on different retrieval tasks. On the Urban1k dataset, direct resize caused a slight decrease in model performance, with Long-CLIP-B and Long-CLIP-L decreasing by 0.9 and 1.5 percentage points on average, respectively. However, on the Chart-to-Text dataset, the same strategy brought significant improvements, with the two models gaining approximately 12.3 and 13.1 percentage points in performance, respectively. This shows that in chart retrieval tasks, retaining the complete structure of the chart (including edge axis labels, legends, etc.) is more important than maintaining precise proportions, proving that our adjustments to chart-specific preprocessing strategies are necessary and effective.
Finding 5: In natural images, a lot of important semantic information is often in the center of the image, so the center crop processing allows the model to pay more attention to the query target. However, in the chart, important information such as the title, x-axis, and y-axis names are all around the chart, which makes it easy for the center crop to cut out important information.
Table 7: Ablation Study on Center Crop. W/ direct resize means we directly resize the image to the required size instead of cropping the chart to the target size.
Models R@10 MRR@5 MRR@10 NDCG@10
Urban1k
Long-CLIP-B 94.1 97.4 85.8 88.6
w/ direct resize 93.2 96.1 85.1 87.8
$\downarrow$0.9 $\downarrow$1.3 $\downarrow$0.7 $\downarrow$0.8
Long-CLIP-L 96.7 98.2 90.7 92.6
w/ direct resize 95.1 97.5 88.9 90.9
$\downarrow$1.6 $\downarrow$0.7 $\downarrow$1.8 $\downarrow$1.5
Chart-to-Text
Long-CLIP-B 87.29 90.38 79.87 82.44
w/ direct resize 98.28 98.85 94.18 95.35
$\uparrow$10.99 $\uparrow$8.47 $\uparrow$14.31 $\uparrow$12.91
Long-CLIP-L 88.51 91.03 81.20 83.61
w/ direct resize 99.35 99.78 97.06 97.75
$\uparrow$10.84 $\uparrow$8.75 $\uparrow$15.86 $\uparrow$14.14
A.5 Exp6: Comparison to Text-to-OCR
Table 8: Performance Comparison of Text-to-Chart Retrieval v.s. Text-to-OCR Retrieval on CRBench Benchmark.
@@ A.5 Exp6: Comparison to Text-to-OCR
R@1 R@5 R@10 MRR@10 NDCG@10 R@1 R@5 R@10 MRR@10 NDCG@10
Text-to-OCR
CLIP 32.79 57.49 65.99 43.27 48.74 9.72 30.36 38.46 18.31 23.12
LongCLIP-B 26.72 57.89 66.80 39.59 46.14 15.38 43.32 57.09 27.70 34.67
LongCLIP-L 23.89 50.20 62.35 35.51 41.91 13.77 44.94 57.49 26.37 33.76
Text-to-Chart
CLIP 12.31 26.15 33.33 19.01 22.43 10.69 21.37 25.95 15.69 18.15
Long-CLIP-B Zhang et al. (2025) 26.67 53.85 62.56 37.99 43.90 22.90 40.46 51.91 30.29 35.33
Long-CLIP-L Zhang et al. (2025) 41.03 75.90 79.49 55.32 55.32 38.93 67.18 74.81 50.99 56.77
ChartFinder 47.18 80.00 84.10 61.23 66.90 41.98 75.57 80.15 55.31 61.40
Experiment Settings: In this experiment, we first utilize PaddleOCR to extract text content from all chart images within the benchmark. We then maintain the original user queries as the retrieval input, while the extracted OCR text serves as the target metadata for each chart. For the implementation of the Text-to-OCR (T2O) paradigm, we directly employ the text encoder of each compared model to encode both the queries and the OCR-derived text, performing retrieval via similarity matching of the resulting embeddings.
Overall Results: Table  8 summarizes the performance of various models under both T2O and Text-to-Chart (T2C) paradigms. The results indicate that for smaller-scale models, T2O performs better in Precise Queries (32.79% R@1) compared to T2C (12.31% R@1), yet its performance in Fuzzy Query settings is remarkably poor (9.72% R@1). As the model parameter scale increases, T2C retrieval demonstrates a significant growth trend, generally surpassing the T2O approach across most overall metrics. Notably, our ChartFinder model achieves the best performance across all tasks and metrics.
Finding 6: For encoders with smaller parameter scales, the T2O paradigm is more effective at processing explicit information by leveraging literal text-matching, as these models have not yet established robust visual perception capabilities. However, as encoder capacity grows, models become more proficient at capturing visual information such as layout, structure, and trends. In these cases, the cross-modal T2C approach demonstrates superior semantic alignment, overtaking pure text retrieval. This confirms that with sufficient capacity, latent visual representations can capture deep analytical insights that remain inaccessible through literal OCR extraction.
Finding 7: Experimental data indicates that analytical intents, such as trend variations and proportional distributions, are typically implicit in the chart’s overall visual layout and cannot be reconstructed from fragmented OCR tokens alone. Consequently, the T2O paradigm cannot satisfy complex retrieval needs. In contrast, cross-modal retrieval is essential for building a semantic bridge between high-level query intent and visual logic, enabling reliable retrieval in fuzzy-intent scenarios.
Appendix B Case Study
@@
Figure 5: Case study on Top-1 Text-to-Chart Retrieval Results of the ChartFinder and CLIP.
@@
(ii) For precise query #1, ChartFinder successfully retrieves the charts that accurately match the query by capturing the keywords and filter conditions urban and rural population correctly. However, the chart CLIP retrieved only reflects the urban population, failing to meet the retrieval requirements of the rural population.
(iiii) For precise query #2, our model aligns the semantic information about the hotel booking channels described by the pie chart with the booking distribution in the query. In contrast, although the retrieved chart of the CLIP is also about hotel information, the description is not about the different ways to book hotels, but the optimization of hotel distribution channels.
(iiiiii) For fuzzy query #3, our model successfully captures the keyword change and retrieves a chart that clearly shows the trend in alcohol consumption. By contrast, although the chart retrieved by CLIP is also related to alcohol consumption, its content only focuses on the absolute ranking of alcohol consumption (“Who’s the Booziest”), completely ignoring the core element of change.
(iviv) For fuzzy query #4, distribution in red refers directly to the concept of data distribution. Our model successfully retrieves a stack area chart showing the distribution of income levels. However, the chart retrieved by CLIP is biased toward the trend of household income growth, although this graph is related to income and does not reflect the analytical information on income level distribution.
Through the four cases above, we can see the effectiveness of our model when matching both precise queries and fuzzy queries with their target charts.
Appendix C Complementary Analysis
C.1 Comparative Analysis with ChartQA
While prior benchmarks such as FigureQA Kahou et al. (2017), DVQA Kafle et al. (2018), and PlotQA Methani et al. (2020) have established robust frameworks for chart understanding, they primarily focus on “single-chart question answering”. In these settings, the model is required to extract graphical primitives, numerical values, or local relations from a specific provided image. ChartInsights Wu et al. (2024) further extends this by evaluating multimodal large language models on low-level analytical perception. However, our proposed text-to-chart retrieval task addresses a distinct and complementary challenge: the discriminative selection of the most appropriate chart from a massive repository containing visually and semantically similar candidates. This paradigm shifts the focus from localized fact extraction within a single chart to the global alignment of natural language intent with latent visual representations across a vast candidate pool. By utilizing our semantic insights synthesis pipeline to bridge the gap between query intent and visual information through multi-level semantic supervision, we provide a new perspective for assessing how models understand the analytical semantics of charts in realistic Business Intelligence (BI) scenarios. In summary, while ChartQA primarily evaluates the comprehension of a single chart in isolation, text-to-chart retrieval demands the simultaneous understanding and comparative analysis of multiple charts to identify the correct target, thereby offering a more rigorous and comprehensive benchmark for the development of real-world retrieval-and-ranking systems.
Table 9: Diagnosis of error-prone word sets in failed retrieval queries.
Error-prone word sets Success rate Failure rate Failure bias
opinion / perception / concern / belief 15.4% 35.9% +20.5%
by / among / across 43.4% 53.2% +9.8%
distribution / percentage / share / proportion / composition / rate 23.1% 19.8% -3.3%
compare / compared / versus (vs) /difference 15.9% 5.9% -10.0%
trend / change / growth / decline / over time / between / from…to 61.0% 47.7% -13.3%
C.2 Error Analysis
To gain a deeper understanding of the model’s limitations, we performed a fine-grained error analysis by conducting a word-level diagnosis on failed queries within CRBench. As summarized in Table 9, retrieval failures are not uniformly distributed but are significantly concentrated in queries involving implicit intent or complex logical structures. Specifically, “opinion-oriented” queries exhibit the highest failure bias of +20.5%, suggesting that the model struggles to align subjective business concerns with concrete visual features when the statistical target is not explicitly stated. Similarly, “multi-dimensional breakdown” queries show a failure bias of +9.8%, highlighting the difficulty in parsing intricate grouping relations where the analytical goal remains implicit .
In contrast, queries with explicit statistical cues such as trend, growth, or ranking yield negative failure biases, indicating that the model effectively captures well-defined analytical patterns with high reliability. These results underscore the necessity of our proposed multi-level semantic supervision to bridge the gap between high-level query intent and latent visual representations.
C.3 The Target-and-Distractor Paradigm
We emphasize that our design is intentionally structured to simulate realistic Business Intelligence (BI) scenarios and to evaluate the fundamental capabilities of retrieval-and-ranking systems.
  • Reason 1: Simulation of Intent-Driven BI Workflows. In practical BI environments, users do not perform 1:1 static mappings; rather, they use natural language to search across a vast repository of existing chart assets. Our task setup mirrors this $1\times N$ search process, shifting the focus from localized fact extraction to the global alignment of natural language intent with visual information across a massive candidate pool.
  • Reason 2: Assessment of Discriminative and Ranking Capabilities. By matching 247 high-quality queries against 21,328 candidates—which include numerous visually and semantically similar distractors—we rigorously assess a model’s ability to distinguish and rank appropriate targets. If a query were generated for every chart, the task would degenerate into a simple image-to-text description matching exercise, failing to evaluate the model’s capacity for comparative analysis among similar candidates.
  • Reason 3: Task Complexity. While benchmarks like Chart-QA evaluate the ability to “read” a single chart in isolation, CRBench addresses the distinct challenge of simultaneous understanding and comparative analysis. Maintaining a large candidate pool relative to the number of queries is essential for developing practical retrieval-and-ranking systems that must operate in high-interference, real-world scenarios.
C.4 Clarification of Paper Contribution
In this section we wish to further clarify the core value of our research across three dimensions—task definition, empirical significance, and methodological depth:
  • Contribution 1: Bridging the Gap Between Academic Focus and Industrial Requirements. While Chart Question Answering (CQA) is widely studied in academia, the challenge of Chart Retrieval—a critical and frequent task in the Business Intelligence (BI) industry—has not received sufficient attention. Our proposed CRBench is the first large-scale benchmark specifically designed for this scenario. It is not merely a dataset but a formal academic modeling of the industrial pain point: “precise target selection from massive chart repositories”.
  • Contribution 2: Revealing Performance Bottlenecks in Realistic BI Scenarios. Through systematic evaluation on CRBench, we provide the first objective evidence that even state-of-the-art multimodal models fail to satisfy industrial-level performance requirements when faced with complex queries and highly similar distractors. This finding highlights a critical shift for the field: moving from “single-chart perception” to “multi-chart discriminative ranking”.
  • Contribution 3: Proposed Data-Driven Pipeline and Scientific Insights. To address these identified performance gaps, we propose a Multi-level Semantic Insights Synthesis Pipeline. Beyond achieving significant performance gains (e.g., nearly doubling results in fuzzy intent scenarios), our exploration has yielded several high-value scientific observations. These include the ··Scaling Law of modality superiority relative to encoder capacity”and the “absolute dependence of fuzzy analytical semantics on latent visual representations”. These insights offer the community a reproducible path for utilizing large-scale synthetic data to solve complex cross-modal alignment problems. In summary, the contribution of this research lies in defining a novel and practical task, revealing an underestimated performance gap, and providing both a feasible solution and a set of foundational scientific insights.
C.5 Data Source Diversity
Figure 6: Domain distribution of data source.
To quantitatively assess the semantic diversity of our collected corpus, we employed a scalable hybrid annotation pipeline on the 8,860 Kaggle CSV datasets. Given the scale of the data, we first utilized an unsupervised learning approach: metadata features (filenames and column headers) were extracted and mapped via TF-IDF, followed by K-Means clustering to group semantically similar datasets. Subsequently, we leveraged GPT-4o to automatically assign domain labels to each cluster by analyzing representative samples. To further resolve the long-tail distribution, we conducted a fine-grained reclassification, successfully identifying distinct sub-domains such as Society and Education.
As shown in Figure 6, the resulting taxonomy spans over 15 distinct domains. While Finance (21.5%) and Media (8.9%) constitute major proportions, the substantial presence of the “Others” category (19.4%) highlights the high-entropy and long-tail nature of our data sources, ensuring the benchmark’s robustness across diverse semantic contexts.
Appendix D Prompt
D.1 Query Generation Prompts for GPT-4o
Query Generation Prompt for GPT-4o You are a user searching for visualizations in a database. You will see 5 visualizations, but you should ONLY focus on the FIRST image when generating queries. The other four images are very similar to the first one, make sure your queries CAN NOT match any of the other four images. Generate two meaningful queries (10-15 words each): 1. Precise Query: Generate a query about specific content in the FIRST visualization that includes: - Meaningful combination of axis labels/categories - Important data values or time ranges - Key categories or measurements - If the image contains time, include the time range in the query. - When covering the time range, do not directly use the initial time range in the chart. - The time range must be within the original chart time span. - Do not use commas in your query. - Do not directly copy text from the image. Use similar wording instead. Example (good): global temperature change 1990-2020 in the United States Bad example: global temperature change 2. Fuzzy Query: Generate a query about the visualization purpose that includes: - Line charts: trend analysis or comparison over time - Bar charts: value comparison or ranking - Pie/stacked charts: distribution or proportion comparison - Scatter plots: correlation or pattern analysis Example (good): annual economic growth comparison between countries Bad example: growth comparison Format your response as: { "Precise query": "10-15 word query about content", "Fuzzy query": "10-15 word query about purpose" } Remember: - ONLY consider the FIRST image - Include relevant context - Be specific and meaningful - Articles and prepositions count as words
D.2 Semantic Insights Generation Prompts for LLaMA
Generation Prompts of Visual-oriented Insights for LLaMA system_prompt = """You are an expert data visualization analyst who excels at crafting clear, engaging narratives. Your task is to write a fluid, well-structured paragraph that describes a data visualization. Write as if explaining the visualization to a professional audience.""" user_prompt = f"""Based on the following chart information, write a single cohesive paragraph explaining the visualization: Title: {chart_data[’describe’]} Subtitle: {chart_data[’sub_title’]} Chart Type: {chart_data[’chart’]} X-axis: {chart_data[’x_name’]} Y-axis: {chart_data[’y_name’]} Categories: {’, ’.join(chart_data.get(’classify’, [])) if chart_data.get(’classify’) else Single category’} Requirements: - Begin with "This {chart_data[’chart’]} chart" - Naturally describe relationships between variables - Use professional yet accessible language - Approximately 100 words """
Generation Prompts of Task-oriented Insights for LLaMA system_prompt = """You are a professional data analyst. Describe the charts purpose and practical applications. Format strictly: Main Purpose: [Single paragraph of 50-100 words describing the visualizations core objective and data presentation approach. Avoid introductory phrases.] """ user_prompt = f"""Chart information: Title: {chart_data[’describe’]} Subtitle: {chart_data[’sub_title’]} Chart Type: {chart_data[’chart’]} X-axis Label: {chart_data[’x_name’]} Y-axis Label: {chart_data[’y_name’]} Data Categories: {’, ’.join(chart_data[’classify’]) if chart_data[’classify’] else None’} """
Generation Prompts of Statistics-oriented Insights for LLaMA system_prompt = """You are an expert data analyst. Based on key statistical metrics, provide analysis. Format strictly: Statistical Analysis: [Single paragraph of 50-100 words analyzing ordering, relationships, ranges, and correlations. Avoid introductory phrases.] """ user_prompt = f"""Chart Information: Title: {chart_data[’describe’]} Chart Type: {chart_type} X-axis: {chart_data[’x_name’]} Y-axis: {chart_data[’y_name’]} Key Statistics: {stats_info} """
Appendix E Ethical and Open Source Statements
E.1 Ethical Statements
Regarding the concern on data compliance, we emphasize that all Kaggle datasets utilized in this work were acquired exclusively through the official Kaggle API. This acquisition process strictly adheres to Kaggle’s Terms of Service and official data access protocols, ensuring that our data collection and utilization are fully compliant with legal requirements and standard research ethics guidelines.
E.2 Open Source Statements
Regarding the inquiry about data availability, we confirm that all research artifacts—including the CRBench benchmark, the fine-tuning datasets, model weights, and the complete training pipeline—will be fully open-sourced upon the acceptance of this paper. We are committed to ensuring transparency and reproducibility to facilitate further research within the community.